# LeanVec: Search your vectors faster by making them fit

### v1

@@ 5 CONCLUSIONS
In this work, we presented LeanVec, a framework that combines linear dimensionality reduction with vector quantization to accelerate similarity search on high-dimensional vectors, including those produced by deep learning models. Additionally, LeanVec speeds up the time-consuming construction of the index used to conduct the search. We presented LeanVec variants for in-distribution (ID) and out-of-distribution (OOD) queries, both leading to state-of-the-art results. LeanVec-OOD uses a novel technique for dimensionality reduction that considers the query and database distributions to simultaneously boost the accuracy and the performance of the framework even further (even matching the performance of LeanVec-ID in the ID setting). Overall, our extensive and varied experiments show that LeanVec yields state-of-the-art results, with an up to 3.7x improvement in search throughput and up to 4.9x faster index build time over the best alternatives.
@@
Appendix A Proof of Proposition 1
@@
Appendix B Proof of Theorem 1
@@
Let $C\geq\max\{C_{f_{{\bm{\mathbf{A}}}}},C_{f_{{\bm{\mathbf{B}}}}}\}$. Summing equations 27 and 28, we get
@@
Finally, plugging equations 34, 35 and 36 in Equation 33 we get
@@
Appendix C Experimental setup.
@@ C.1 Baseline approaches
C.2 Metrics
@@ Appendix D Datasets
D.1 New text-to-image dataset with OOD queries
We introduce wit-512-1M, a new dataset with OOD queries stemming from a text-to-image application. The WIT dataset [^9] https://github.com/google-research-datasets/wit is a multimodal multilingual dataset that contains 37 million rich image-text examples extracted from Wikipedia pages. For each example in the first million, we take the image [^10] Images downloaded from https://storage.cloud.google.com/wikimedia-image-caption-public/image_data_train.tar and encode it using the multimodal OpenAI CLIP-ViT-B32 model (Radford et al. 2021) to generate a database vector. We create the query set using the first $2\cdot 10^{4}$ text descriptions in one of the provided test sets [^11] Downloaded from https://storage.googleapis.com/gresearch/wit/wit_v1.test.all-00000-of-00005.tsv.gz (concatenating the Reference and Attribution description fields) and generating the corresponding embeddings using CLIP-ViT-B32-multilingual-v1  (Reimers & Gurevych 2020). [^12] The use of CLIP-ViT-B32 for images and multi-lingual CLIP-ViT-B32-multilingual-v1 for text follows the protocol suggested in https://huggingface.co/sentence-transformers/clip-ViT-B-32-multilingual-v1. Finally, for each query, we compute the 100 ground truth nearest neighbors using maximum inner product. We use the first $10^{4}$ queries as a learning set and the remaining $10^{4}$ as a test set.
D.2 New question-answering dataset with ID and OOD queries
@@
Figure 12: Search throughput and accuracy comparison for different LenVec combinations. We observe no noticeable differences in the graphs built with and without dimensionality reduction.

### v2

@@ 5 CONCLUSIONS
In this work, we presented LeanVec, a framework that combines linear dimensionality reduction with vector quantization to accelerate similarity search on high-dimensional vectors, including those produced by deep learning models. Additionally, LeanVec speeds up the time-consuming construction of the index used to conduct the search. We presented LeanVec variants for in-distribution (ID) and out-of-distribution (OOD) queries, both leading to state-of-the-art results. LeanVec-OOD uses two novel techniques for dimensionality reduction that consider the query and database distributions to simultaneously boost the accuracy and the performance of the framework even further (even matching the performance of LeanVec-ID in the ID setting). Overall, our extensive and varied experiments show that LeanVec yields state-of-the-art results, with an up to 3.7x improvement in search throughput and up to 4.9x faster index build time over the best alternatives.
@@
Appendix A Speeding up graph construction with dimensionality reduction
In the introduction, we discuss the bottlenecks observed when searching with high-dimensional vectors using a graph index. However, these difficulties extend to the construction process of the graph index itself. In every graph index, the construction process can be divided into two main steps: search and pruning.
We start from a directed graph $G=({\mathcal{X}},E)$, where the database vector set ${\mathcal{X}}$ is used as the node set and the edge set $E$ is initialized depending on the specific graph-construction algorithm (Malkov & Yashunin 2018, Jayaram Subramanya et al. 2019), where we may even start with $E=\emptyset$. In order to keep the search complexity bounded, each node in the graph has a maximum out-degree $R$. To build the graph, we iteratively perform the following update routine for each node ${\bm{\mathbf{x}}}\in{\mathcal{X}}$:
Search
We first run the search algorithm using the node ${\bm{\mathbf{x}}}$ as the query on the current graph $G$, seeking a set of ${\mathcal{C}}$ of approximate nearest neighbors with cardinality larger than $R$.
Pruning
We use ${\mathcal{C}}$ as a set of candidate nodes to form outgoing edges (or arcs) from ${\bm{\mathbf{x}}}$. To increase the navigability of the graph, a pruning algorithm (Arya & Mount 1993, Malkov & Yashunin 2018, Jayaram Subramanya et al. 2019) is run on ${\mathcal{C}}$, yielding a set ${\mathcal{C}}^{\prime}\subseteq{\mathcal{C}}$ such that $|{\mathcal{C}}^{\prime}|<R$. We then replace all the arcs in $E$ starting from ${\bm{\mathbf{x}}}$ with the set $\{({\bm{\mathbf{x}}},{\bm{\mathbf{x}}}^{\prime})\,|\,{\bm{\mathbf{x}}}^{\prime}\in{\mathcal{C}}^{\prime}\}$. It is important to note that all pruning algorithms rely on computing distances between pairs of vectors in ${\bm{\mathbf{C}}}$.
Any slowdowns caused by working with high-dimensional vectors will carry over directly to the graph construction process. The runtime of the search and pruning algorithms are dominated by fetching high-dimensional vectors from memory and computing distances on them. Zhang et al. (2022) observed that graph construction can be accelerated by reducing the vector dimensionality. However, because the dimensionality reduction technique proposed by Zhang et al. (2022) has a time-consuming inference algorithm (based on transformers), it cannot be applied for search: its runtime exceeds the runtime of the search itself). LeanVec applies equally to the search and graph construction processes by alleviating memory pressure while remaining computationally lean.
Scalability. The graph construction technique detailed above is executed (at least once) for each node in the graph (i.e., for each vector in the database). Thus, the technique scales linearly with the graph size both in the number $n$ of nodes and in the number of edges (this quantity is upper bounded by $nR$). Consequently, the LeanVec acceleration has a linear impact on the graph construction runtime.
Appendix B Proof of Proposition 1
@@
Appendix C Convergence analysis of the Frank-Wolfe algorithm for LeanVec-OOD
We consider the general constrained problem
$\min_{z\in{\mathcal{D}}}f(z),$ (22)
We now present two extensions of $\|\nabla f(z^{(t)})\|$ and the Lipschitz assumption, which are standard in unconstrained optimization, suitable for constrained optimization (Lacoste-Julien 2016).
Definition 1.
The Frank-Wolfe gap at $z^{(t)}$ is defined as
Definition 2.
Lemma 1 (Jaggi (2013, Lemma 7)).
[…]
In our case, $z=\{{\bm{\mathbf{A}}},{\bm{\mathbf{B}}}\}$, $f$ is defined in Equation 10, and ${\mathcal{D}}=\{z\,|\,{\bm{\mathbf{A}}}\in{\mathcal{C}}\land{\bm{\mathbf{B}}}\in{\mathcal{C}}\}$ for ${\mathcal{C}}$ defined in Equation 2. ${\mathcal{D}}$ is convex, being the intersection of two convex sets, and $f$ is a non-convex function of $z$. The partial derivatives in Equation 13 are linear, and thus L-Lipschitz continuous, and $\operatorname{diam}_{\|\cdot\|}({\mathcal{C}})$ is bounded. This implies that the curvature for ${\bm{\mathbf{A}}}$ (resp. ${\bm{\mathbf{B}}}$), given a fixed ${\bm{\mathbf{B}}}$ (resp. ${\bm{\mathbf{A}}}$) is finite.
We are now ready to state our main convergence result.
Theorem 1.
Proof.
@@ Appendix C Convergence analysis of the Frank-Wolfe algorithm for LeanVec-OOD
Let $C\geq\max\{C_{f_{{\bm{\mathbf{A}}}}},C_{f_{{\bm{\mathbf{B}}}}}\}$. Summing equations 29 and 30, we get
@@
Finally, plugging equations 36, 37 and 38 in Equation 35 we get
@@
Interestingly, our setting is very related to one in the recent work by Peng & Vidal (2023). They study Equation 22 with a BCD algorithm, but concentrate on the exact minimization of the subproblems or in a case where retractions are applied. We leave as future work the extension of Theorem 1 and its proof to the general setting with more than two blocks.
Appendix D Experimental setup.
@@ D.1 Baseline approaches
We also considered RoarANN (Chen et al. 2024), the winner of the OOD track of the NeurIPS’23 Big-ANN competition (Simhadri et al. 2024), using the hyperparameters used by its authors for the competition. RoarANN failed with a segfault while building indices for open-images-512-13M and rqa-768-10M.
D.2 Finding the optimal target dimensionality $d$
In our experiments, we only vary one parameter to tune the performance: the target dimensionality $d$. The hyperparameters used to build the indices are set once and shared for all datasets (described in the introductory paragraph of Appendix D). The choice of LVQ-8 for the primary vectors and either LVQ-8 or FP16 for the secondary vectors is consistently superior to the other choices (see Figure 10). Thus, we also keep this choice fixed throughout the experiments.
The target dimensionality $d$ that yields the best performance is dataset dependent as the loss we are optimizing depends on the data distribution (see the ablation study in Figure 9). However, for most datasets whose dimensionality is typical of modern deep learning models, there is an optimum that lies somewhere in the range [160, 256]. Notice that the differences between the achieved performance when $d$ is in this range are not striking. Any of these choices would yield significant improvements over the state of the art. Even a poor choice, like $d=320$ is in many cases, already provides significant performance gains over the state of the art. A user not seeking to do any hyperparameter tuning will be covered by picking $d\in[160,256]$.
We point out that tuning $d$ is no different than tuning, for example, the number of segments used in a product quantizer, the number of levels using in a residual quantizer, or the number of clusters used in a traditional vector quantizer. With any vector compression method, its optimal behavior will depend on finding the best fit to a specific data distribution. It is important, though, that its behavior degrades gracefully as we move away from the best-fit hyperparameters. Our experiments show that LeanVec does exhibit such a graceful degradation.
For each dataset, we use the dimensionality $d$ that yields the highest search performance at 90% accuracy (see LABEL:table:datasets). To find this $d$, we use the same graph built from uncompressed vectors and measure the search throughput by increasing $d$ in steps of 32 starting from 64. This involves building projection matrices, projecting the database vectors for each $d$, and finally running searches to find the best $d$.
D.3 Metrics
@@ Appendix E Datasets
E.1 New text-to-image dataset with OOD queries
We introduce wit-512-1M, a new dataset with OOD queries stemming from a text-to-image application. The WIT dataset [^9] https://github.com/google-research-datasets/wit is a multimodal multilingual dataset that contains 37 million rich image-text examples extracted from Wikipedia pages. For each example in the first million, we take the image [^10] Images downloaded from https://storage.cloud.google.com/wikimedia-image-caption-public/image_data_train.tar and encode it using the multimodal OpenAI CLIP-ViT-B32 model (Radford et al. 2021) to generate a database vector. We create the query set using the first $2\cdot 10^{4}$ text descriptions in one of the provided test sets [^11] Downloaded from https://storage.googleapis.com/gresearch/wit/wit_v1.test.all-00000-of-00005.tsv.gz (concatenating the Reference and Attribution description fields) and generating the corresponding embeddings using CLIP-ViT-B32-multilingual-v1  (Reimers & Gurevych 2020). [^12] The use of CLIP-ViT-B32 for images and multi-lingual CLIP-ViT-B32-multilingual-v1 for text follows the protocol suggested in https://huggingface.co/sentence-transformers/clip-ViT-B-32-multilingual-v1. Finally, for each query, we compute the 100 ground truth nearest neighbors using maximum inner product. We use the first $10^{4}$ queries as a test set and the remaining $10^{4}$ as a learning set.
E.2 New question-answering dataset with ID and OOD queries
@@
Figure 13: Search throughput and accuracy comparison between the two LeanVec-OOD variants, namely LeanVec-FW (LABEL:algo:leanvec) and LeanVec-ES (LABEL:algo:leanvec_eig_search).
Figure 14: Search throughput and accuracy comparison for different LeanVec combinations. We observe no noticeable differences in the graphs built with and without dimensionality reduction.
Relative errors
${\color[rgb]{0.12109375,0.46484375,0.70703125}\definecolor[named]{pgfstrokecolor}{rgb}{0.12109375,0.46484375,0.70703125}\left\|n_{s}^{-1}{\bm{\mathbf{Q}}}_{n_{s}}{{\bm{\mathbf{Q}}}_{n_{s}}}^{\top}-m^{-1}{\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top}\right\|_{F}/\left\|m^{-1}{\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top}\right\|_{F}}$
${\color[rgb]{1,0.49609375,0.0546875}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.49609375,0.0546875}\left\|n_{s}^{-1}{\bm{\mathbf{X}}}_{n_{s}}{{\bm{\mathbf{X}}}_{n_{s}}}^{\top}-n^{-1}{\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}\right\|_{F}/\left\|n^{-1}{\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}\right\|_{F}}$
${\color[rgb]{0.171875,0.62890625,0.171875}\definecolor[named]{pgfstrokecolor}{rgb}{0.171875,0.62890625,0.171875}\left\|{{\bm{\mathbf{A}}}_{n_{s}}}^{\top}{\bm{\mathbf{A}}}_{n_{s}}-{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{A}}}\right\|_{F}/\left\|{{\bm{\mathbf{A}}}}^{\top}{\bm{\mathbf{A}}}\right\|_{F}}$
${\color[rgb]{0.83984375,0.15234375,0.15625}\definecolor[named]{pgfstrokecolor}{rgb}{0.83984375,0.15234375,0.15625}\left\|{{\bm{\mathbf{B}}}_{n_{s}}}^{\top}{\bm{\mathbf{B}}}_{n_{s}}-{{\bm{\mathbf{B}}}}^{\top}{\bm{\mathbf{B}}}\right\|_{F}/\left\|{{\bm{\mathbf{B}}}}^{\top}{\bm{\mathbf{B}}}\right\|_{F}}$
${\color[rgb]{0.58203125,0.40234375,0.7421875}\definecolor[named]{pgfstrokecolor}{rgb}{0.58203125,0.40234375,0.7421875}\left\|{{\bm{\mathbf{P}}}_{n_{s}}}^{\top}{\bm{\mathbf{P}}}_{n_{s}}-{{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}\right\|_{F}/\left\|{{\bm{\mathbf{P}}}}^{\top}{\bm{\mathbf{P}}}\right\|_{F}}$
Figure 15: The error between a sample covariance matrix and its expectation converges quickly (Koltchinskii & Lounici 2017) with a growing sample size (at a $\sqrt{n}$ rate). This quick convergence carries over to the LeanVec-OOD loss being a function of ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}={\bm{\mathbf{Q}}}{{\bm{\mathbf{Q}}}}^{\top}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}={\bm{\mathbf{X}}}{{\bm{\mathbf{X}}}}^{\top}$. In this experiment, we first optimize the LeanVec-OOD loss, using the full learning sets with $n=10^{6}$ database vectors and $m=10^{4}$ queries, with LABEL:algo:leanvec to obtain ${\bm{\mathbf{A}}}$ and ${\bm{\mathbf{B}}}$ or with LABEL:algo:leanvec_eig_search to obtain ${\bm{\mathbf{P}}}$. Then, using different numbers $n_{s}$ of random subsamples to compute ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$, we obtain ${\bm{\mathbf{A}}}_{n_{s}}$, ${\bm{\mathbf{B}}}_{n_{s}}$, and ${\bm{\mathbf{P}}}_{n_{s}}$ using the corresponding algorithm. We can see that the relative errors (formulas on the right-hand side) drop quickly in the left plot as $n_{s}$ grows.
Figure 16: The brute-force search accuracy of LeanVec-ES is robust to subsampling the set of query vectors used for learning/training. Let $n_{s}$ be this query sample size. Some degradation can be perceived when using $n_{s}=D$ or $n_{s}=2D$ samples but it vanishes when using $n_{s}=4D$ samples or more. This result is in agreement with the quick convergence of the matrices ${\bm{\mathbf{K}}}_{{\bm{\mathbf{Q}}}}$ and ${\bm{\mathbf{K}}}_{{\bm{\mathbf{X}}}}$ defined in Equation 8 as $n_{s}$ grows.
Refer to captionRefer to captionRefer to captionRefer to caption
rqa-768-1M
laion-512-1M
wit-512-1M
t2i-200-1M
Figure 17: The loss in Equation 14 is a smooth function of $\beta$ when ${\bm{\mathbf{P}}}=\operatorname{eigsearch}(\beta)$ and empirically we observe that it has a unique minimizer (different for each $d$). LABEL:algo:leanvec_eig_search finds the minimum (marked with a circle) of this loss using a derivative-free scalar minimization. We point out that the solutions with $\beta=0$ and $\beta=1$ correspond to the SVD of the query and database vectors, respectively.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 18: LeanVec-FW, LeanVec-ES, and SVD correspond to LABEL:algo:leanvec, to LABEL:algo:leanvec_eig_search, and to the algorithm in Section 2.1, respectively. LeanVec-ES+FW corresponds to initializating LABEL:algo:leanvec with the output of LABEL:algo:leanvec_eig_search. Although LeanVec-ES often yields a lower loss value than LeanVec-FW, their brute-force search performance are comparable. The behaviors of LeanVec-ES and LeanVec-ES+FW are almost indistinguishable across the board, bringing assurance about the good empirical performance of LeanVec-ES.